MP3 report – week 15 6/4/2010 – 11/04/2010 9 interventions + 10A/s ramp rate testing - PowerPoint PPT Presentation

mp3 report week 15 6 4 2010 11 04 2010 9 interventions 10a s ramp rate testing n.
Download
Skip this Video
Loading SlideShow in 5 Seconds..
MP3 report – week 15 6/4/2010 – 11/04/2010 9 interventions + 10A/s ramp rate testing PowerPoint Presentation
Download Presentation
MP3 report – week 15 6/4/2010 – 11/04/2010 9 interventions + 10A/s ramp rate testing

play fullscreen
1 / 7
MP3 report – week 15 6/4/2010 – 11/04/2010 9 interventions + 10A/s ramp rate testing
135 Views
Download Presentation
tessa
Download Presentation

MP3 report – week 15 6/4/2010 – 11/04/2010 9 interventions + 10A/s ramp rate testing

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. MP3 report – week 15 6/4/2010 – 11/04/20109 interventions + 10A/s ramp rate testing MP3/QPS team in CCC Monday Morning Meeting : 12th of April 2010

  2. Extraction switches problems • Intervention on Monday for a RB.A78 energy extraction switch. • Micro-switch problem • Intervention on Tuesday for a RQTD.A81B1 energy extraction switch • No real problem. Procedures updated Monday Morning Meeting : 12th of April 2010

  3. Undulator trip on Tuesday. • Trip of the RU.L4 ; slow drift of U_res ; can be included in combo rampdown-precycle and precyclesubsequences when main circuits are at the lowest current Monday Morning Meeting : 12th of April 2010

  4. RQTD/RQTF trips from feedback • Feedback incompatible with QPS protection • Meeting on Friday organized by A. Siemko • A solution seems possible by changing the time discrimination. Monday Morning Meeting : 12th of April 2010

  5. Ramp rate fix type test on Thursday • Type test performed on S56; new cards installed for the EE with a delay of 400ms in the odd points and 580ms in the even points. Monday Morning Meeting : 12th of April 2010

  6. Ramp rate fix type test • Successive tests with FPA via PIC at 2, 4 , 6 kA while ramping up to 10A/s  no magnet triggered! • Test successful, reverted to the old cards for operation (at 2A/s ramp rate). Monday Morning Meeting : 12th of April 2010

  7. MP3 issues since 6th of April 2010 • The second part of the week was remarkably quiet with very few problems • RCBXH1.R1: power cycle problem: in case of need for a power cycle, access is needed. Will be fixed in next technical stop • RQT13.L8B1: operational limit of 300A is not enough to squeeze in LHCb – limit changed to 352A. Medium term plan .... Monday Morning Meeting : 12th of April 2010